OpenStudy (smurfy14):

simplify (-32a^10)^(3/5)

OpenStudy (anonymous):

First, it's important to recognize that 32 and \(a^{10}\) are 5th powers. So, taking fifth roots, we get \((-2a^{2})^{3}\). Are you good from here?

OpenStudy (smurfy14):

no. what do you mean 5th roots??

OpenStudy (anonymous):

So 5th roots are kind of like square roots

OpenStudy (anonymous):

4 is the square of 2, so \(\sqrt{4}=2\).

OpenStudy (anonymous):

\(2^{5}=32\), so \(32^{1/5}=2\).

OpenStudy (smurfy14):

ok and you get a^2 by dividing 10 by 5 right?...so what happened to the denominator of five outside the ()?

OpenStudy (anonymous):

That goes away because we divided the powers by 5.

OpenStudy (smurfy14):

...ok so is that the final answer?

OpenStudy (anonymous):

Well no, now you still have to cube it.

OpenStudy (smurfy14):

-8a^6 correct?

OpenStudy (anonymous):

Perfect.

OpenStudy (smurfy14):

ok cool so how would you do (9^(1/2)+16^(1/2))^-2?

OpenStudy (anonymous):

Ick, that's an ugly one. Are you sure there is a plus in there?

OpenStudy (smurfy14):

haha ya

OpenStudy (anonymous):

Stupid me, it's way easy, 9 and 16 are squares.

OpenStudy (smurfy14):

oh ok i see it :P do you know what 7^-2 would be?

OpenStudy (anonymous):

\[\frac{1}{7^{2}}\], right?, which then becomes? (leave it in a fraction)

OpenStudy (smurfy14):

oh ok! if you dont mind could you help me with like four more? i have this as a precalc summer assignment and we havent learned any of this yet.

OpenStudy (anonymous):

Actually, I can't sorry, I have to take off. I am tutoring someone at 5.

OpenStudy (smurfy14):

ok well thanks for the help!

OpenStudy (anonymous):

You're very welcome, sorry I couldn't help you on the other four.

OpenStudy (smurfy14):

its fine.